# The Failed Migration of Academic Twitter

### v1

@@ 3.4 Twitter and Mastodon cross-platform analysis
To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 2). We plot the daily frequency of these tweets over time in Figure 3. Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement.
@@
"Looks like Mastodon isn’t an option anyway. Academic servers aren’t taking new accounts at the moment."
"I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am <scholar’s Mastodon account>."
"As you know, I’m over on Mastodon too. Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does."
@@
Figure 3: Dynamic frequency of Twitter conversations related to academics migrating to Mastodon.
@@
"Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity. but the *big* servers that most ppl are on are… kinda academic? kinda techy? very… not fun? it feels like linkedin or facebook groups to a lot of people."
"My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity."
@@
"Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd"
@@
Amongst the discussions, we observe concurrent mentions of other platforms such as Bluesky [^3] https://bsky.app/ and Threads [^4] https://www.threads.net/. These have diluted the migration as users explore various platforms for different purposes. For instance,
@@
"I’m curious re categories of people I follow and where they go. IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?)."
@@
"Hey if anyone’s still here on science twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic"
3.4.1 Case Study: Physical Sciences and Mathematics Scholars.
We observe that scholars in the physical sciences and mathematics were amongst the most engaged on Mastodon. We randomly sampled 100 scholars in these fields who had provided Twitter account information and matched their Mastodon accounts with their Twitter accounts. We recorded whether they are still active on Twitter, and whether they still provide their Mastodon account information in their username, profile description, or pinned tweets. Additionally, we collected follower and following statistics of the sampled scholars on Twitter, all statistics shown in Table 3. We observe that 86% of sampled scholars are still active on Twitter, and among these active users, 62.8% do not provide Mastodon information in easily accessible places, e.g., username, bio, pinned tweets, indicating no signs of migration or intention to migrate. These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform. The rest of the active users still make their Mastodon information accessible to the community. For instance, one scholar noted:
"Hi ya’ll for those moving over to #Mastodon you can currently find me <Scholar’s Mastodon Account>."
This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon. In contrast, 5% of the sampled scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter. For example, one scholar’s last post reads,
"I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!"
Some, 9% of our sampled scholars, chose to lock their accounts as a sign of leaving Twitter, leaving a message in their profile description directing followers to their Mastodon accounts.
Contain Mastodon Info Not Contain Mastodon Info Following Count Follower Count
Active 32 54 Mean: 1593, std: 2684 Mean:11022, std: 30146
Inactive 4 1 Mean: 330, std: 451 Mean: 848, std: 1396
Locked 7 2 Mean: 1551, std:1361 Mean: 1631, std: 1837
Table 3: Engagement statistics of 100 sampled scholars.
In addition to segmentation, many scholars direct their followers to Threads and Bluesky instead of Mastodon, with 2% providing Threads information and 6% providing Bluesky information among the sampled scholars. This trend indicates that Mastodon faces significant competition from other emerging platforms, making it even more challenging for Mastodon to become a preferred alternative for academic discourse.
We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4. Our analysis reveals that 55% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users). Among those who are inactive on Twitter or have locked their accounts, 71% are long-term or persistent users of Mastodon. This indicates that while Mastodon has some loyal users, their numbers are relatively small compared to the entire academic community.
One_Time Short_Term Long_Term Persistent
Active 12 19 28 27
Inactive 1 1 2 1
protected 0 2 1 6
Table 4: Mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
4 Discussion and Conclusion
Our analysis of professional communities migrating between social media platforms highlights the unique challenges these users face during such transitions. While decentralized platform offered a promising alternative, the established communities and clear audience engagement on Twitter proved too significant to overcome. The fragmented communities on Mastodon lacked the cohesiveness found on Twitter. Our results show that the initial surge in Mastodon sign-ups did not translate into sustained user engagement.
Among the users we tracked, the majority failed to maintain the same level of activity. This could be attributed to steep learning curve compared to twitter as well as competition from more user-friendly platforms like Bluesky and Threads. These alternatives might have drawn away potential users exacerbating Mastodonś difficulties in becoming a central hub for professionals communities. Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter which is crucial for sustaining professional discourse.
@@ 4 Discussion and Conclusion
Our study highlights that without a good alternative, professionals will naturally gravitate back to familiar platforms where their communities are already established through years. Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement. While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions. Successful transitioning involves creating communities that rival established social media networks and user experiences of platforms such as Twitter.
{credits}
[…]
There is no competing interest.

### v2

@@ 4.4. Analysis of Tweets on Mastodon Migration
To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 3). We plot the daily frequency of these tweets and engagement counts over time aggregated by week in Figure 8. Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement. After the migration period, although conversations were less frequent, they received more attention (likes), indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
@@
”I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am ¡scholar’s Mastodon account¿.”
”As you know, I’m over on Mastodon too. Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does.”
@@
Figure 8. Dynamic frequency of Twitter conversations related to academics migrating to Mastodon and corresponding engagement metrics (connection and like counts), aggregated by week.
@@
”Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity. but the *big* servers that most ppl are on are… kinda academic? kinda techy? very… not fun? it feels like linkedin or facebook groups to a lot of people.”
”My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.”
@@
”Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd”
@@
Amongst the discussions, we again observe concurrent mentions of other platforms such as Bluesky and Threads. These have diluted the migration as users explore various platforms for different purposes. For instance,
@@
”I’m curious re categories of people I follow and where they go. IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?).”
@@
”Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic”
From a qualitative perspective, these conversations reflect both optimism in Mastodon’s potential and the practical concerns academics face during the migration.
5. Discussion and Conclusion
In this work, we explored the migratory dynamics of Academic Twitter communities to Mastodon. Through a series of analyses, we highlight that building a community on a new platform requires more than a coordinated effort and availability of alternatives.
Results from longitudinal analysis of Mastodon user data showed that the initial surge in sign-ups did not translate into sustained long-term user engagement. Most recent user activity on Twitter reveals that many academics in our dataset returned to using Twitter actively. Further analysis of engagement metrics reveals that the level of established history, as well as the strong communities established on Twitter, with some over a decade, proved too significant to overcome. Without a complete migration of user history, and established community, professionals will naturally gravitate back to familiar platforms where their communities are already established through years.
Several factors likely contributed to the failure of Mastodon to retain users. First, its decentralized structure comes with a steep learning curve compared to Twitter. Furthermore, competition from more user-friendly platforms like Bluesky and Threads potentially drew users away from Mastodon. Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter. While some users expressed doubts, certain groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition.
One significant limitation of this study is the inconsistency in data collection timing across different platforms. This discrepancy could potentially introduce biases in the analysis, as activity levels of scholars might vary depending on when the data was captured. Additionally, incorrect or incomplete social media handles provided by the scholars led to gaps in data that we could not manually resolve due to the scale of the dataset. Moreover, the changes to platform access (such as Twitter API rate limitations) restricted our ability to gather sufficient data throughout the study. However, we believe our findings from the remaining data is substantial and provides valuable insights for the cross platform analyses.
@@ 5. Discussion and Conclusion
Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement. While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions. However, the Academic Twitter community’s large scale attempt to migrate from Twitter demonstrates their adaptability and resilience. If any group could successfully make this migration, it would be a community like Academic Twitter, united by a shared purpose and strong networks. Yet, as encouraging as these efforts are, there is a long way to go.